# Inner products of deep level Deligne–Lusztig representations of Coxeter type

### v1

@@ Abstract.
In this article we prove orthogonality relations for deep level Deligne–Lusztig varieties of Coxeter type, attached to a reductive group over a local non-archimedean field, which splits over an unramified extension. This extends results of [DI24], where the quasi-split case was handled. This allows to construct new irreducible representations of parahoric subgroups of $p$-adic groups.
@@ 1. Introduction
Let $G$ be a reductive group over $k$, which splits over $\breve{k}$. Let $T\subseteq B$ be a maximal torus and a Borel subgroup of $G$, such that $T$ splits and $B$ becomes rational over $\breve{k}$. Denote by $U$ the unipotent radical of $B$ and assume that $(T,U)$ is a Coxeter pair (see §2.1). In particular, $T$ is elliptic and the apartment of $T$ in the reduced Bruhat–Tits building of $G$ consists of one point. Bruhat–Tits theory attaches to this point a (connected) parahoric $\mathcal{O}_{k}$-model $\mathcal{G}$ of $G$. For any $r\leq\infty$ we can regard $\mathcal{G}(\mathcal{O}_{{\breve{k}}}/\varpi^{r})=\mathbb{G}_{r}(\overline{}\mathbb{F}_{q})$ as the geometric points of a perfect ¯$\overline{}\mathbb{F}_{q}$-scheme $\mathbb{G}_{r}$. This is done via the (truncated, if $r<\infty$) positive loop functor, see e.g. [Zhu17, §1.1] (or [DI24, §2]) for details.
We fix now some $r<\infty$ and write $\mathbb{G}=\mathbb{G}_{r}$. If $\mathcal{G}$ is defined over $\mathcal{O}_{k}$, then $\mathbb{G}$ is defined over $\mathbb{F}_{q}$; in this case we also denote by $F$ the geometric Frobenius of $\mathbb{G}$, so that $\mathbb{G}^{F}=\mathbb{G}(\mathbb{F}_{q})$. Moreover, if $H\subseteq G$ is a subscheme, then we denote by $\mathcal{H}$ its closure in $\mathcal{G}$ and by $\mathbb{H}\subseteq\mathbb{G}$ the corresponding subscheme of $\mathbb{G}$. Consider the closed subscheme
(1.1) $X=X^{\mathcal{G}}_{T,U,r}=\{x\in\mathbb{G}\colon x^{-1}F(x)\in F\mathbb{U}\}$
of $\mathbb{G}$. The group $\mathbb{G}^{F}\times\mathbb{T}^{F}$ acts on it by $(g,t)\colon xgxt$. Let $\ell\neq p$ be a prime. For any smooth character $\chi\colon\mathbb{T}^{F}\rightarrow\overline{}\mathbb{Q}_{\ell}^{\times}$ we have the virtual $\mathbb{G}^{F}$-module
$H_{c}^{\ast}(X)[\chi]=\sum_{i\in\mathbb{Z}}(-1)^{i}H^{i}(X_{T,U},\overline{}\mathbb{Q}_{\ell}^{\times})[\chi],$
where for any $\mathbb{T}^{F}$-module $M$, $M[\chi]$ denotes the $\chi$-isotypic subspace. The following is our main result.
@@
Recently, under a mild condition on $p$, Chan [Cha24] shows that the inner product formula in Theorem 1.1 holds if $(T,\theta)$ is split-generic, using a different approach.
Now we comment on Theorem 1.1. First, we explain why “it suffices” to establish the theorem for a single Coxeter pair $(T,U)$. Ultimately, we are interested in smooth representations of the $p$-adic group $G(k)$. One has the $p$-adic Deligne–Lusztig space $X_{T,U}\cong\dot{X}_{\dot{c}}(b)$ from [Iva23a, §7.2 and §11.2] equipped with an action of $G(k)\times T(k)$. By [Nie24, Iva23b], one has
$X_{T,U}=\coprod_{\gamma\in G(k)/\mathcal{G}(\mathcal{O}_{k})}\gamma X_{T,U,\infty}^{\mathcal{G}}.$
Thus for any smooth character $\theta\colon T(k)\rightarrow\overline{}\mathbb{Q}_{\ell}^{\times}$, the $\theta$-isotypic part of the cohomology of $X_{T,U}$ equals
$R_{T,U}(\theta)={\text{c-}\rm Ind}_{\mathcal{G}(\mathcal{O}_{k})Z(k)}^{G(k)}H^{\ast}_{c}(X)[\theta|_{\mathcal{T}(\mathcal{O}_{k})}],$
@@
where $Z\subseteq G$ is the center and $H^{\ast}_{c}(X)[\theta|_{\mathcal{T}(\mathcal{O}_{k})}]$ is extended to a representation of $\mathcal{G}(\mathcal{O}_{k})Z(k)$ in the unique way such that its central character is $\theta|_{Z(k)}$. Now, the point is that by [Iva23a, Corollary 7.25, Lemma 11.3], $X_{T,U}$ are all mutually $G(k)\times T(k)$-equivariantly isomorphic, when $(T,U)$ varies through all Coxeter pairs $(T,U)$ with a fixed $T$. Thus, the $G(k)$-representation $R_{T,U}(\theta)$ is independent of the choice of $U$, as long as $(T,U)$ remains Coxeter. So, it suffices to know the statement of the theorem for at least one Coxeter pair. In fact, our proof shows that for many groups $G$ Theorem 1.1 holds for all pairs $(T,U)$, see Remark 2.4.
@@
Next, we explain why the condition on $q$ in the theorem is very mild, so that Theorem 1.1 even allows to construct new irreducible representations. Recall that by the work of Yu and Kaletha [Yu01, Kal19], one can attach a supercuspidal irreducible $G(k)$-representation $\pi_{(S,\theta)}$ to any regular elliptic pair $(S,\theta)$ consisting of a maximal elliptic torus $S\subseteq G$ and a sufficiently nice smooth character $\theta\colon S(k)\rightarrow\overline{}\mathbb{Q}_{\ell}^{\times}$. A crucial point for this to work is the existence of a Howe factorization of $\theta$, cf. [Kal19, §3.6]. However, not all characters admit a Howe factorization, when the residue characteristic $p$ is small and $G$ is not an inner form of $\mathrm{GL}_{n}$.
@@
For instance, if $p\in\{2,3,5\}$, there exist many examples of pairs $(T,\theta)$ with $T$ unramified Coxeter (hence covered by our main result when $q$ satisfied condition (2.1) – in particular, whenever $q>5$) such that ${\rm Stab}_{W_{e}^{F}}(\theta)=\{1\}$, but $\theta$ does not admit a Howe factorization. For examples of $(T,\theta)$ not admitting a Howe factorization we refer to the forthcoming work of Fintzen–Schwein [FS], where an algebraic approach to the extension of Yu’s construction is pursued. As ${\rm Stab}_{W_{e}^{F}}(\theta)=\{1\}$, one should expect an irreducible supercuspidal $G(k)$-representation attached to $(T,\theta)$, but Yu’s construction does not apply as there is no Howe factorization. The point is now that our cohomological construction does not require any condition on $p$, but only a mild one on $q$. In particular, there are many examples of $k,G,T,\theta$ such that ±$\pm H^{\ast}_{c}(X)[\theta|_{\mathcal{T}(\mathcal{O}_{k})}]$ is an irreducible $\mathcal{G}(\mathcal{O}_{k})$-representation, which does not appear in Yu’s construction.
However, let us also note that our method is not yet complete, in the sense that it does not prove that the resulting $G(k)$-representation $R_{T,U}(\theta)$ is irreducible supercuspidal. So far there is no purely Deligne–Lusztig theoretic proof of this fact; the closest purely Deligne–Lusztig theoretic result states (in the case of inner forms of $\mathrm{GL}_{n}$) that $R_{T,U}(\theta)$ is admissible and hence a direct sum of finitely many irreducible supercuspidal representations, see [CI23, Theorem 6.1].
@@ 2.1. More notation
We use the notation from the introduction. Moreover, we denote by $Z$ the center of $G$, $N_{G}(T)$ the normalizer of $T_{\breve{k}}$ in $G_{\breve{k}}$, by $W=N_{G}(T)/T$ the Weyl group of $T$, by $X^{\ast}(T)$ (resp. $X_{\ast}(T)$) the group of characters (resp. cocharacters) of $T_{\breve{k}}$ and by $\langle\cdot,\cdot\rangle\colon X^{\ast}(T)\times X_{\ast}(T)\mathbb{Z}$ the natural pairing. We write $\Phi$ for the root system of $T_{\breve{k}}$ in $G_{\breve{k}}$, $\Phi^{+}$ for the subset of positive roots determined by $B$, and $\Delta\subseteq\Phi^{+}$ for the subset of positive simple roots. We write $S\subseteq W$ for the corresponding set of simple reflections.
Let $c\in W$ be the unique element such that $FB={}^{c}B$. Then for any lift $\dot{c}$ of $c$, ${\mathrm{Ad}}({\dot{c}})^{-1}\circ F:G(\breve{k})G(\breve{k})$ fixes the pinning $(T,B)$, hence defines automorphisms, denoted by $\sigma$, of the based root system $\Delta\subseteq\Phi$ and of the Coxeter system $(W,S)$. Note that $\sigma$ does not depend on the choice of the lift $\dot{c}$. We call $(T,B)$ (or $(T,U)$) a Coxeter pair if $c$ is a $\sigma$-Coxeter element in the Coxeter triple $(W,S,\sigma)$, that is, if a(ny) reduced expression of $c$ contains precisely one element from each $\sigma$-orbit on $S$. Moreover, we assume throughout the article that $c$ is $\sigma$-Coxeter, and hence $(T,U)$ is a Coxeter pair.
Except for $G$, $\mathcal{G}$ and their subgroups (which are defined over $k,\breve{k}$ resp. $\mathcal{O}_{k},\mathcal{O}_{\breve{k}}$), all schemes appearing below are perfect schemes perfectly of finite presentation and perfectly smooth over ¯$\overline{}\mathbb{F}_{q}$. For a review of perfect geometry we refer to [Zhu17, Appendix A]. We freely make use of the $6$-functor formalism of étale cohomology for such schemes with $\overline{}\mathbb{Q}_{\ell}$-coefficients. Moreover, for a perfect ¯$\overline{}\mathbb{F}_{q}$-scheme we denote by $H_{\ast}(Y)=H_{c}^{\ast}(Y,\overline{}\mathbb{Q}_{\ell})$ its $\ell$-adic étale cohomology with compact support.
@@ 2.3. A condition on $q$
Let $\omega_{\alpha}$ denotes the fundamental coweight of $\alpha\in\Delta$. For a $\sigma$-orbit $\mathcal{O}\subseteq\Delta$ of simple roots, we set $\omega_{\mathcal{O}}^{\vee}=\sum_{\alpha\in\mathcal{O}}\omega_{\alpha}^{\vee}$, where $\omega_{\alpha}^{\vee}$ denotes the fundamental coweight of $\alpha\in\Delta$. We prove our main result under the following condition on $q$:
@@ 2.5. Support
Then one checks that $(c\sigma)^{-i_{0}}=\sigma^{-i_{0}}w$ for some $w\in W$ such that $\operatorname{supp}(w)\in S-\{s_{\alpha}\}$. Hence $\alpha\in\operatorname{supp}(w(\gamma))$ and $\alpha_{j}=\sigma^{-i_{0}}(\alpha)\in\operatorname{supp}(\sigma^{-i_{0}}w(\gamma))=\operatorname{supp}((c\sigma)^{-i_{0}}(\gamma))$. So we can assume that $\alpha=\alpha_{j}$. Let $10\leqslant i\leqslant\delta-$ 1. Note that $(c\sigma)^{i}=u_{i}\sigma^{i}$ for some $u_{i}\in W$ with $\operatorname{supp}(u_{i})\in S-\{\sigma^{i}(\alpha_{j})\}$. It follows that $\sigma^{i}(\alpha)\in\operatorname{supp}((c\sigma)^{i}(\gamma))$. So the statement follows. ∎
@@ 3.1. The schemes $\Sigma^{i}$
$\Sigma^{i}=\{(x,x^{\prime},y)\in\mathbb{U}\times F^{i+1}\mathbb{U}\times\mathbb{G};xF(y)=yx^{\prime}\}.$
@@
$\dim_{\overline{}\mathbb{Q}_{\ell}}H_{\ast}(\Sigma_{w}^{0})=\begin{cases}H_{\ast}((\dot{w}\mathbb{T})^{c\sigma})&\text{if $w\in W_{e}^{c\sigma}$,}\\ 0&\text{otherwise.}\end{cases}$

### v2

# Deep level Deligne–Lusztig representations of Coxeter type

@@ Abstract.
In this article we study the cohomology of deep level Deligne–Lusztig varieties of Coxeter type, attached to a reductive group over a local non-archimedean field, which splits over an unramified extension. This allows to construct some new irreducible representations of parahoric subgroups of $p$-adic groups. Moreover, in the quasi-split case we prove that these compactly induce to finite direct sums of irreducible supercuspidal representations of the $p$-adic group. This extends previous results of [DI24], [CI23].
@@ 1. Introduction
Let $G$ be a reductive group over $k$, which splits over $\breve{k}$. Let $T\subseteq B$ be a maximal torus and a Borel subgroup of $G$, such that $T$ splits and $B$ becomes rational over $\breve{k}$. Denote by $W$ the Weyl group of $T$ in $G$. Denote by $U$ resp. $U^{-}$ the unipotent radicals of $B$ resp. the opposite Borel subgroup and assume that $(T,U)$ is a Coxeter pair (see §2.1). Attached to $(T,U)$ there is a $p$-adic Deligne–Lusztig space, on geometric points given by
$X_{T,U}=\{g\in G(\breve{k})\colon g^{-1}F(g)\in(U^{-}\cap FU)(\breve{k})\}.$
It admits a continuous action of $G(k)\times T(k)$ given by $(g,t)\colon xgxt$. See [Iva23a, §7 and §11] (and §5.2 below). Let $\theta\colon T(k)\overline{}\mathbb{Q}_{\ell}^{\times}$ be a smooth character. The $\theta$-isotypic component $R_{T}^{G}(\theta)$ of the homology of $X_{T,U}$ is an object in the (derived) category of smooth $G(k)$-representations, cf. [IM]. The goal of this article is to further investigate properties of $R_{T}^{G}(\theta)$, extending and generalizing results from [DI24, CI23].
To describe our results we need more notation. The apartment of $T$ in the reduced Bruhat–Tits building of $G$ consists of one point. Bruhat–Tits theory attaches to this point a (connected) parahoric $\mathcal{O}_{k}$-model $\mathcal{G}$ of $G$. By [Iva23b, Nie24], $X_{T,U}$ admits a decomposition $X_{T,U}=\coprod_{\gamma\in G(k)/\mathcal{G}(\mathcal{O}_{k})}\gamma X_{T,U}^{\mathcal{G}}$, where
$X_{T,U}^{\mathcal{G}}=\{g\in\mathcal{G}(\mathcal{O}_{\breve{k}})\colon g^{-1}F(g)\in(\mathcal{U}^{-}\cap F\mathcal{U})(\mathcal{O}_{\breve{k}})\}$
is an affine ¯$\overline{}\mathbb{F}_{q}$-scheme (here $\mathcal{U}$ denotes the closure of $U$ in $\mathcal{G}$).
Fix some $r\leq\infty$. We can regard $\mathcal{G}(\mathcal{O}_{{\breve{k}}}/\varpi^{r})=\mathbb{G}(\overline{}\mathbb{F}_{q})$ as the geometric points of a perfect $\mathbb{F}_{q}$-scheme $\mathbb{G}=\mathbb{G}_{r}$. This is done via the (truncated, if $r<\infty$) positive loop functor, see e.g. [Zhu17, §1.1] (or [DI24, §2]) for details. For a subscheme $H\subseteq G$, we denote by $\mathcal{H}$ its closure in $\mathcal{G}$ and by $\mathbb{H}\subseteq\mathbb{G}$ the corresponding subscheme of $\mathbb{G}$. We denote by $F$ the geometric Frobenius of $\mathbb{G}$, so that $\mathbb{G}^{F}=\mathbb{G}(\mathbb{F}_{q})$. Then $X_{T,U}^{\mathcal{G}}$ is isomorphic to the inverse limit over $r$ of its truncations in each $\mathbb{G}_{r}$. Each of these truncations is a perfectly smooth perfect ¯$\overline{}\mathbb{F}_{q}$-scheme, and up to an $\mathbb{A}^{n}$-bundle (not affecting the cohomology), it equals
(1.1) $X=X^{\mathcal{G}}_{T,U,r}=\{x\in\mathbb{G}\colon x^{-1}F(x)\in F\mathbb{U}\},$
Note that $X$ is equipped with the action of the finite group $\mathbb{G}^{F}\times\mathbb{T}^{F}$ given by $(g,t)\colon xgxt$.
By these geometric considerations ($+\varepsilon$), $R_{T}^{G}(\theta)$ admits the following more explicit description (which might, for the purposes of this article, also be considered as a definition). Let $Z\subseteq G$ denote the center of $G$. For any $\mathbb{T}^{F}$-module $M$, let $M[\chi]$ denote the $\chi$-isotypic subspace. Then, if $\theta|_{\mathcal{T}(\mathcal{O}_{k})}$ factors through a character $\chi$ of $\mathbb{T}^{F}$, then
$R_{T}^{G}(\theta)={\rm cInd}_{\mathcal{G}(\mathcal{O}_{k})Z(k)}^{G(k)}H_{c}^{\ast}(X)[\chi],$
where $H_{c}^{\ast}(X)[\theta]=\sum_{i\in\mathbb{Z}}H_{c}^{i}(X,\overline{}\mathbb{Q}_{\ell})[\theta]$ is the $\ell$-adic equivariant Euler characteristic of $X$ (regarded as a virtual $\mathbb{G}^{F}$-module), inflated to a $\mathcal{G}(\mathcal{O}_{k})$-representation, and extended to $\mathcal{G}(\mathcal{O}_{k})Z(k)$ in the unique way such that $Z(k)$ acts by $\theta|_{Z(k)}$. Our first main result concerns the representations in the cohomology of $X$.
@@
Recently, under a mild condition on $p$, Chan [Cha24] shows by a different approach that the inner product formula holds in a much more general case, which in particular includes the case that $T$ is elliptic.
Our second main result concerns the cuspidality of the compactly induced $G(k)$-representation $R_{T}^{G}(\theta)$. It generalizes [CI23, Theorem 6.1].
Theorem 1.3.
Assume that $G$ is unramified and that $q$ satisfies condition (2.1). Let $\theta\colon T(k)\overline{}\mathbb{Q}_{\ell}^{\times}$ be smooth with trivial stabilizer in $W^{F}$. Then $R_{T}^{G}(\theta)$ is up to sign a finite direct sum of irreducible supercuspidal representations of $G(k)$.
@@
Some comments on our results are in order. First, we explain why “it suffices” to establish Theorem 1.1 for a single Coxeter pair $(T,U)$. Ultimately, we are interested in the smooth $G(k)$-representation $R_{T}^{G}(\theta)$. By [Iva23a, Corollary 7.25, Lemma 11.3], $X_{T,U}$ are mutually $G(k)\times T(k)$-equivariantly isomorphic, when $(T,U)$ varies through all Coxeter pairs $(T,U)$ with a fixed $T$ [^1] This is not clear for the schemes $X_{T,U,r}$ at least if $G$ is not quasi-split (for the quasi-split case, see [DI24, Corollary 4.1.4]). Thus, $R_{T}^{G}(\theta)$ is independent of the choice of $U$. So, it suffices to know the statement of Theorem 1.1 for at least one Coxeter pair. In fact, our proof shows that for many groups $G$ Theorem 1.1 holds for all pairs $(T,U)$, see Remark 2.4.
@@
Next, we explain why the condition on $q$ in Theorems 1.1 and 1.3 is very mild, so that the theorems even gives rise to new supercuspidal representations of $G(k)$. Recall that by the work of Yu and Kaletha [Yu01, Kal19], one can attach a supercuspidal irreducible $G(k)$-representation $\pi_{(S,\theta)}$ to any regular elliptic pair $(S,\theta)$ consisting of a maximal elliptic torus $S\subseteq G$ and a sufficiently nice smooth character $\theta\colon S(k)\rightarrow\overline{}\mathbb{Q}_{\ell}^{\times}$. A crucial point for this to work is the existence of a Howe factorization of $\theta$, cf. [Kal19, §3.6]. However, not all characters admit a Howe factorization, when the residue characteristic $p$ is small and $G$ is not an inner form of $\mathrm{GL}_{n}$.
@@
For instance, if $p\in\{2,3,5\}$, there exist many examples of pairs $(T,\theta)$ with $T$ unramified Coxeter (hence covered by Theorem 1.1 when $q$ satisfied condition (2.1) – in particular, whenever $q>5$) such that ${\rm Stab}_{W_{e}^{F}}(\theta)=\{1\}$, but $\theta$ does not admit a Howe factorization. For examples of $(T,\theta)$ not admitting a Howe factorization we refer to the forthcoming work of Fintzen–Schwein [FS], where an algebraic approach to the extension of Yu’s construction is pursued. As mentioned in [CO23], since ${\rm Stab}_{W_{e}^{F}}(\theta)=\{1\}$ one should expect an irreducible supercuspidal $G(k)$-representation attached to $(T,\theta)$, but Yu’s construction does not apply as there is no Howe factorization. The point is now that our cohomological construction does not require any condition on $p$, but only a mild one on $q$. In particular, there are many examples of $k,G,T,\theta$ such that ±$\pm H^{\ast}_{c}(X)[\theta|_{\mathcal{T}(\mathcal{O}_{k})}]$ is an irreducible $\mathcal{G}(\mathcal{O}_{k})$-representation, which does not appear in Yu’s construction. Moreover, Theorem 1.3 implies that its induction to $G(k)$ is supercuspidal.
@@ 2.1. More notation
We use the notation from the introduction. Moreover, we denote by $N_{G}(T)$ the normalizer of $T_{\breve{k}}$ in $G_{\breve{k}}$, so that $W=N_{G}(T)/T$ is the Weyl group of $T$, by $X^{\ast}(T)$ (resp. $X_{\ast}(T)$) the group of characters (resp. cocharacters) of $T_{\breve{k}}$ and by $\langle\cdot,\cdot\rangle\colon X^{\ast}(T)\times X_{\ast}(T)\mathbb{Z}$ the natural pairing. We write $\Phi$ for the root system of $T_{\breve{k}}$ in $G_{\breve{k}}$, $\Phi^{+}$ for the subset of positive roots determined by $B$, and $\Delta\subseteq\Phi^{+}$ for the subset of positive simple roots. We write $S\subseteq W$ for the corresponding set of simple reflections.
Let $c\in W$ be the unique element such that $FB={}^{c}B$. Then for any lift $\dot{c}$ of $c$, ${\mathrm{Ad}}({\dot{c}})^{-1}\circ F:G(\breve{k})G(\breve{k})$ fixes the pinning $(T,B)$, hence defines automorphisms, denoted by $\sigma$, of the based root system $\Delta\subseteq\Phi$ and of the Coxeter system $(W,S)$. Note that $\sigma$ does not depend on the choice of the lift $\dot{c}$. We call $(T,B)$ (or $(T,U)$) a Coxeter pair if $c$ is a $\sigma$-Coxeter element in the Coxeter triple $(W,S,\sigma)$, that is, if a(ny) reduced expression of $c$ contains precisely one element from each $\sigma$-orbit on $S$. Moreover, we assume until the end of §4 that $c$ is $\sigma$-Coxeter, and hence $(T,U)$ is a Coxeter pair.
Except for $G$, $\mathcal{G}$ and their subgroups (which are defined over $k,\breve{k}$ resp. $\mathcal{O}_{k},\mathcal{O}_{\breve{k}}$), all schemes appearing below are perfect schemes perfectly of finite presentation and perfectly smooth over ¯$\overline{}\mathbb{F}_{q}$. For a review of perfect geometry we refer to [Zhu17, Appendix A]. We freely make use of the $6$-functor formalism of étale cohomology for such schemes with $\overline{}\mathbb{Q}_{\ell}$-coefficients. Moreover, we fix a prime number $\ell\neq p$, and for a perfect ¯$\overline{}\mathbb{F}_{q}$-scheme we denote by $H^{\ast}(Y)=H_{c}^{\ast}(Y,\overline{}\mathbb{Q}_{\ell})$ its $\ell$-adic étale cohomology with compact support.
@@ 2.3. A condition on $q$
Let $\omega_{\alpha}^{\vee}$ denotes the fundamental coweight of $\alpha\in\Delta$. For a $\sigma$-orbit $\mathcal{O}\subseteq\Delta$ of simple roots, we set $\omega_{\mathcal{O}}^{\vee}=\sum_{\alpha\in\mathcal{O}}\omega_{\alpha}^{\vee}$, where $\omega_{\alpha}^{\vee}$ denotes the fundamental coweight of $\alpha\in\Delta$. We prove our main result under the following condition on $q$:
@@ 2.5. Support
Then one checks that $(c\sigma)^{-i_{0}}=\sigma^{-i_{0}}w$ for some $w\in W$ such that $\operatorname{supp}(w)\subseteq\Delta-\{\alpha\}$. Hence $\alpha\in\operatorname{supp}(w(\gamma))$ and $\alpha_{j}=\sigma^{-i_{0}}(\alpha)\in\operatorname{supp}(\sigma^{-i_{0}}w(\gamma))=\operatorname{supp}((c\sigma)^{-i_{0}}(\gamma))$. So we can assume that $\alpha=\alpha_{j}$. Let $10\leqslant i\leqslant\delta-$ 1. Note that $(c\sigma)^{i}=u_{i}\sigma^{i}$ for some $u_{i}\in W$ with $\operatorname{supp}(u_{i})\subseteq\Delta-\{\sigma^{i}(\alpha_{j})\}$. It follows that $\sigma^{i}(\alpha)\in\operatorname{supp}((c\sigma)^{i}(\gamma))$. So the statement follows. ∎
@@ 3.1. The schemes $\Sigma^{i}$
$\Sigma^{i}=\{(x,x^{\prime},y)\in F\mathbb{U}\times F^{i+1}\mathbb{U}\times\mathbb{G};xF(y)=yx^{\prime}\}.$
@@
$H_{\ast}(\Sigma_{w}^{0})=\begin{cases}H_{\ast}((\dot{w}\mathbb{T})^{c\sigma})&\text{if $w\in W_{e}^{c\sigma}$,}\\ \{0\}&\text{otherwise.}\end{cases}$